OpenStudy (anonymous):

Simplify: -27a³b² / 3ab A. –9a²b B. –9a³b² C.–3ab D. -a²/9b

OpenStudy (anonymous):

http://oi61.tinypic.com/66ixbp.jpg

OpenStudy (anonymous):

divide \[\frac{ a ^{3} }{ a }\frac{ b ^{2} }{ b }\frac{ -27 }{ 3 }\]

OpenStudy (anonymous):

then recombine

OpenStudy (muzzack):

B